### v1

# Stochastic Path Planning in Correlated Obstacle Fields

Abstract
We introduce the Stochastic Correlated Obstacle Scene (SCOS) problem, a navigation setting with spatially correlated obstacles of uncertain blockage status, realistically constrained sensors that provide noisy readings and costly disambiguation. Modeling the spatial correlation with Gaussian Random Field (GRF), we develop Bayesian belief updates that refine blockage probabilities, and use the posteriors to reduce search space for efficiency. To find the optimal traversal policy, we propose a novel two-stage learning framework. An offline phase learns a robust base policy via optimistic policy iteration, augmented with information bonus to encourage exploration in informative regions, followed by an online rollout policy with periodic base updates via a Bayesian mechanism for information adaptation. This framework supports both Monte Carlo point estimation and distributional reinforcement learning (RL) to learn full cost distributions, leading to stronger uncertainty quantification. We establish theoretical benefits of correlation-aware updating and convergence property under posterior sampling. Comprehensive empirical evaluations across varying obstacle densities, sensor capabilities demonstrate consistent performance gains over baselines. This framework addresses navigation challenges in environments with adversarial interruptions or clustered natural hazards.
00footnotetext: This work was supported by ONR Grant N00014-22-1-2572 and NSF Award #2319157.
Keywords: Stochastic navigation, sequential decision making, Bayesian update, distributional reinforcement learning, Gaussian Random Fields, mutual information
1 Introduction
Navigation and path planning in uncertain, partially observable environments presents a critical challenge in operations research and robotics, particularly when complicated by additional factors including adversarial interruptions and environmental obstacles. Applications span network interdiction problems with incomplete information (Azizi and Seifi, (2024), Smith and Song, (2020)), autonomous driving (Wang et al.,, 2024), disaster response and fire evacuation (Shiri and Salman,, 2019), and adversarial route planning in military operations (Berger et al., (2012), Hickling et al., (2023)). In such contexts, obstacles may have two statuses: false obstacles correspond to false alarms, while true obstacles represent actual threats requiring costly detours or clearance. Classic models, building upon foundational frameworks such as the Canadian Traveler’s Problem (CTP) (Bar-Noy and Schieber,, 1991) and the Stochastic Obstacle Scene (SOS) problem (Papadimitriou and Yannakakis,, 1991) typically simplify the problem by making strong assumptions rarely met in practice: (i) obstacle blockages are independent (Alkaya et al.,, 2021), ignoring potential spatial patterns arising from adversarial behavior or terrain features (Alkaya et al.,, 2021; Pitilakis et al.,, 2016), and (ii) sensing is often idealized, assuming noiseless reveals with no explicit sensing cost or global sensing with high accuracy (Lamarre and Kelly,, 2025), whereas real-world sensors have limited detection range and varying reliability levels.
To address these limitations, we introduce the Stochastic Correlated Obstacle Scene (SCOS) model, which extends the SOS framework by explicitly incorporating correlation among obstacles and realistic sensors with limited range and varying accuracy level. In this setting, a navigating agent moves through an environment containing spatially correlated disk-shaped obstacles (e.g., danger zones) whose blockage nature is revealed only through noisy sensor readings within a limited sensing range and costly disambiguation upon contact, introducing a strategic tradeoff between information collection costs and traversal risks. These enhancements make the SCOS model particularly suitable for real-world challenges, including navigation through natural disaster zones or defense operations in adversarial context.
The SCOS presents two major decision-making challenges: (i) the agent must optimally balance exploration, gathering information from highly uncertain paths to potentially reduce future traversal cost, and exploitation, taking a low-risk path with lower immediate expected cost based on current knowledge. (ii) The problem requires effective handling of correlation information. Since sensor readings and disambiguation outcomes not only reveal local blockage information but also provide insights about correlated, unobserved regions. These challenges are difficult to address using existing approaches (Eyerich et al.,, 2010; Lim et al.,, 2017; Blumenthal and Shani,, 2023; Lamarre and Kelly,, 2025), which exhibit various limitations including limited robustness under high uncertainty, lack of theoretical performance guarantees or frameworks for continuous belief refinement under spatial correlations.
To tackle these challenges, we propose a two-stage policy learning strategy with the following key contributions: (i) Exploration guided by information gain: we augment the learning process with information bonus built upon mutual information, encouraging deeper exploration of uncertain and informative regions compared to traditional methods, with convergence property established. (ii) Flexible value function estimate approaches: in additional to classic Monte Carlo methods providing point estimates, our framework supports distribution RL to learn full cost distributions, providing stronger uncertainty quantification and more robust policies under high noise and risk. (iii) Offline-online integration: we develop a two-stage strategy enabling online adaptation, which outperforms standard rollout methods relying on static base policies. It consists of an offline phase that learns a high quality base policy via hybrid RL, and an online rollout phase with periodic base updates via a Bayesian mechanism that improves point estimates and distribution approximation. (iv) Belief updates with space reduction: blockage probabilities are refined using a Bayesian updating framework based on Gaussian Random Field model, with a search space reduction step guided by the posteriors to enhance computational efficiency.
The article proceeds as follows: Section 2 reviews related work and Section 3 formally defines the SCOS problem and introduce its formulation as a sequential decision model. Section 4 describes the Bayesian updating process for obstacle information based on Gaussian Random Field (GRF). Section 5 introduces three crucial building blocks of our proposed two-stage policy learning framework, followed by detailed descriptions of the offline learning strategy and online rollout policy with base updates in Sections 6 and 7. Section 8 presents empirical evaluations of our approach.
2 Related Work
Stochastic path planning problems have been extensively studied, with the Canadian Traveler Problem (CTP) (Bar-Noy and Schieber,, 1991) representing a key graph-theoretical foundation and providing insights into navigating graphs with incrementally revealed edge statuses. The Stochastic Obstacle Scene (SOS) problem (Papadimitriou and Yannakakis,, 1991) extends these insights to continuous obstacles settings, emphasizing dynamic learning and information gathering cost. However, simplified assumptions like obstacle independence and unrestricted sensor capability have limited the practical applicability of SOS and existing variants (Aksakalli et al., (2011), Aksakalli and Ari, (2014), Alkaya and Oz, (2017), Alkaya et al., (2021)).
Network Interdiction Problem (Smith and Song, (2020), Azizi and Seifi, (2024)) shares conceptual similarities with SOS and CTP, but usually focuses on edges interdiction from an adversary’s (i.e., interdictor) perspective with assumed perfect information. By contrast, our problem provides a complementary aspect. We prioritize path planning from the traveler’s perspective, managing uncertainty, dynamic exploration costs and spatially correlated obstacles causing realistic area blockage effects.
Spatial correlation in navigation has been explored under the CTP framework. The Gaussian Traveler Problem (GTP) (Dey et al.,, 2014) models correlations between edge travel time or costs using a Gaussian process (GP), but assumes all edges are traversable, limiting its applicability to the setting with blockages where costly disambiguation is required. Similarly, the risk-averse CVaR-CTP (Lamarre and Kelly,, 2025) allows correlated edge costs and updates from noiseless observations. The Bayesian Canadian Traveler Problem (BCTP) (Lim et al.,, 2017; Hou and Srinivasa,, 2022) incorporates correlated edge blockages, but relies on the strong assumption of a restrictive prior hypothesis space. In contrast, we use Gaussian Random Field (GRF) model to represent spatial correlations, as a generalization of Gaussian processes for structured spatial domains. This model allows us to model continuous spatial correlations among obstacles without restricting the hypothesis space and supports a complete posterior updating under noisy sensing.
To manage exploration-exploitation tradeoffs in uncertain environments, deterministic threshold-based policies were proposed (Koenig and Likhachev, (2002), Bnaya et al., (2009), Lim et al., (2017)), which lack a clear method to determine an appropriate threshold value and overlook the probabilistic information. Policies based on rollout (Eyerich et al., (2010), Hou and Srinivasa, (2022), Blumenthal and Shani, (2023)) and UCT (Upper Confidence Bounds Applied to Trees) related policies (Kocsis and Szepesvári, (2006), Tolpin and Shimony, (2012)) improve decision-making by using probabilistic sampling and trajectory simulation. However, they are sensitive to the quality of simulation policies and the accuracy level of probabilistic information, with no guarantee of convergence to optimality. Penalty-based policies modify traversal costs by adding penalties to discourage high-risk paths (Alkaya et al., (2015), Sahin and Aksakalli, (2015), Alkaya et al., (2021)), offering computationally efficient but requiring manual hyperparameter tuning with performance heavily dependent on problem settings.
Reinforcement learning (RL) approaches (Sutton and Barto,, 2018) have been successfully applied to balance immediate and long-term gains in navigation problems (Yu and Bertsekas,, 2013; Polydoros and Nalpantidis,, 2017; Wang et al.,, 2020). RL methods can broadly be categorized into model-based and model-free approaches. Model-based approaches construct a model of environment’s dynamics for efficient exploration and fast convergence, but suffer from high computational requirements and sensitivity to model accuracy. Model-free approaches bypass the need for explicit modeling, but their reliance on extensive exploration often leads to slower convergence. Hence, hybrid approaches combining the strength of both have emerged, shown to achieve promising performance (Silver et al., (2008), Feinberg et al., (2018), Pinosky et al., (2023)).
We build upon these foundations, unifying their strength while addressing limitations, through a framework that integrates correlation modeling, efficient exploration, and real-time decision-making, as detailed in the following sections.
3 The Stochastic Correlated Obstacle Scene Problem (SCOS)
The Stochastic Correlated Obstacle Scene (SCOS) extends the original SOS framework by modeling correlated obstacles and realistic sensor limitations. The SCOS models an agent traversing from a starting point to a goal point within a traversal region containing disk-shaped obstacles whose blockage statuses are initially unknown.
Figure 1: SCOS environment with obstacle statuses (left, red and gray disks representing true and false obstacles, respectively) versus with probability estimates in sensing range (right).
Formally, let the two-dimensional traversal region be $\Omega\subset\mathbb{R}^{2}$, with a set of disk-shaped obstacles located at $X=X^{F}\cup X^{O}$, where $X^{F}$ and $X^{O}$ represent locations of false and true obstacles, respectively. False obstacles can be traversed through, while true obstacles block traversal. Each obstacle $x\in X$ has a fixed $\text{radius}(x)>0$. An undirected graph $\mathcal{G}$ is imposed over $\Omega$, consisting of a set of vertices $\mathcal{V}(\mathcal{G})$ representing discrete navigation locations, and $\mathcal{E}(\mathcal{G})$ represents a set of feasible undirected edges along which the agent navigates between vertices. For each edge $e\in\mathcal{E}(\mathcal{G})$, undirected graph implies that $e=(v_{i},v_{j})$ and $e=(v_{j},v_{i})$ define the same edge, where $v_{i},v_{j}\in\mathcal{V}(\mathcal{G})$.
The agent is equipped with a noisy sensor providing local blockage probability estimates, $\tilde{\rho}_{x}\in[0,1]$, of each obstacle $x$ lying within certain range $R$. These probability estimates can be updated when obstacles are re-encountered, enabling dynamic belief revision. Figure 1 illustrates this scenario, showing the contrast between actual obstacle states and the agent’s probabilistic beliefs within its sensing range. Nearby obstacles exhibit spatial correlations, motivated by strategic placements or natural clustering. These dependencies are captured via Gaussian Random Field (GRF), where blockage probabilities are inferred from a spatial correlated latent process. Additionally, the agent can disambiguate each obstacle’s status with certainty once arriving at a vertex with an edge adjacent to the obstacle, incurring a disambiguation cost $c(x)$.
The agent can only traverse edges that do not intersect true or ambiguous obstacles that have not been disambiguated. When encountering an ambiguous obstacle, the agent performs costly disambiguation, then either traverses through the obstacle if it is false, or takes a detour around the blockage if it is true. Consequently, the total traversal cost from starting to goal point is a random variable that depends on the blockage status of encountered obstacles.
Given a starting point designated as a specific vertex, denoted as $s$, and a goal point denoted as $g$, the objective is formulated as:
$\min_{p\in\mathcal{P}_{sg}}\mathbf{E}\left[L_{p}+C_{p}\right],$
where $\mathcal{P}_{sg}$ denotes the set of all paths from $s$ to $g$ in $\mathcal{G}$, $L_{p}$ represents the random variable denoting the traversal length of a path $p$, and $C_{p}$ is the random variable indicating the disambiguation costs associated with traversing path $p$.
3.1 Formulating SCOS as a Sequential Decision Problem
Because disambiguations and sensor estimates continuously provide new information about the traversal region, SCOS is naturally posed as a sequential decision problem. At each decision step, the agent collects new information and updates the belief about the traversal region, then chooses the subsequent traversal path and the disambiguation location. Previous works typically formulate such problems as deterministic Partially Observable Markov Decision Process (DET-POMDP) (Bonet,, 2012; Dey et al.,, 2014; Aksakalli et al.,, 2016). However, solving DET-POMDPs is PSPACE-complete (Bonet,, 2012), and the model transitions between successive steps are often complicated. To better facilitate the decision process, we instead model the problem using a universal framework with belief state proposed for sequential decision problems (Powell,, 2019).
Adapted to our problem setting, this framework includes five components at each discrete time step $t\in\{0,1,...,T\}$, where a decision is required:
  • State Variables, $S_{t}=\{\mathcal{V}_{t},B_{t}\}$, contains information needed for decision making. $\mathcal{V}_{t}$ represents the agent’s physical location (i.e., the vertex in the discretized graph). $B_{t}$ is the belief state capturing the agent’s knowledge about obstacle uncertainty. $B_{t}=\{X_{t}^{O},X_{t}^{F},X_{t}^{U}\}$, where $X_{t}^{O}$, $X_{t}^{F}$, and $X_{t}^{U}$ denote the sets of true obstacles, false obstacles and ambiguous obstacles, respectively. For ambiguous obstacles, their blockage probabilities are included, denoted as $\rho(X_{t}^{U})=\{\rho_{x}:x\in X_{t}^{U}\}$. We use $s_{t}$, $v_{t}$ and $b_{t}$ for their specific realizations at $t$.
  • Decision, $d_{t}$, represents the agent’s decision of the next obstacle-free path segment to follow. Rather than selecting an action at every graph edge, we let the agent choose a macro–path: an obstacle-free path segment that terminates either (i) at the first vertex adjacent to an ambiguous obstacle, or (ii) at the goal $g$. This approach reduces the decision complexity from potentially hundreds of edge level choices to only a few macro-decisions. Then, the decision $d_{t}$ can be viewed as selecting the stopping vertex, since the path segment is uniquely determined by the current location and the chosen stopping point. We can also denote it as $d(s_{t})$ to emphasize its dependency on state variables.
  • Exogenous Information, $W_{t}$, includes new information available to the agent. Specifically, $W_{t+1}$ contains the actual status of obstacles observed after performing $d_{t}$, as well as the new sensor estimates for obstacles within the sensing range. It can be empty if the agent reaches $g$ directly without any sensing or disambiguation.
  • Transition Function, $S_{t+1}=\mathcal{T}(S_{t},d_{t},W_{t+1})$, describes the state information evolution based on the current state variables $S_{t}$, decision $d_{t}$ and new observation $W_{t+1}$. The new physical location $V_{t+1}$ becomes the vertex chosen by $d_{t}$, while the belief state $B_{t+1}=\{X^{O}_{t+1},X^{F}_{t+1},X^{U}_{t+1}\}$ is updated based on the new obstacle knowledge provided in $W_{t+1}$. The transition uncertainty arises from the stochastic nature of obstacle blockage. For $x\in X^{U}_{t+1}$, blockage probabilities are updated using a Bayesian mechanism incorporating spatial correlations between obstacles. This transition process is depicted in Figure 2.
  • Objective Function seeks an optimal policy $\pi^{*}$ that minimizes the expected total traversal cost. The immediate cost $\mathcal{C}(S_{t},d_{t})$ includes the Euclidean length $\ell_{t}$ of traversing to the stopping vertex, and the disambiguation cost $c_{t}$ if required. Therefore, the objective function is formulated as:
    $\pi^{*}=\operatorname*{arg\,min}_{\pi}\mathbf{E}\left(\sum_{t=0}^{T}\mathcal{C}(S_{t},d^{\pi}(S_{t}))\right),$
    where $T$ denotes the (random) arrival time, $\mathcal{C}(S_{t},d^{\pi}(S_{t}))=\ell_{t}+c_{t}$, and $d^{\pi}(S_{t})$ is the decision taken at $S_{t}$ under policy $\pi$.
$S_{t}$$d_{t}$$W_{t+1}$$S_{t+1}$decisionobservationnext state
Figure 2: A schematic illustration of transitions in the decision-making process
Table A1 in Appendix summarizes the key notations used for problem formulation and belief representation.
4 Updating Framework
Due to the inter-dependency (i.e., correlation) between obstacles’ blockage probabilities, sensor readings and disambiguation outcomes at one location provide useful information about the status of all ambiguous obstacles. This motivates an online update mechanism that iteratively refines beliefs, allowing decisions to be made based on increasingly accurate information. We establish the theoretical benefit to justify the proposed correlation-aware updating approach in Section 4.4.
Previous work on Gaussian classification for binary object status updates model correlations through a latent variable, followed by a squash step to transform values into probability scale (Kapoor et al.,, 2010). However, the resulting likelihood function complicates the derivation of posterior and predictive distributions. Approximation techniques (Nickisch et al.,, 2008) can mitigate this issue, but introduce additional computational cost and concerns on approximation accuracy, especially when used in sequential updates.
To address these challenges, we map each obstacle’s blockage probability $\rho_{x}\in(0,1)$ to log-odds values $\log\left(\frac{\rho_{x}}{1-\rho_{x}}\right)$. This transformation is a common practice in occupancy mapping problems (O’Callaghan and Ramos, (2012), Li et al., (2018)). We then develop a Bayesian updating framework using the Gaussian Random field (GRF) to model the spatial structure of these log-odds values. This strategy enables exact posterior derivation for sequential updates.
4.1 Prior Information and Assumptions
Following the problem definition in Section 3, the traversal region contains $n$ obstacles located at positions $X=X^{O}\cup X^{F}=\{x_{1},x_{2},\ldots,x_{n}\}$. Each obstacle has an unknown binary status denoted as $Z=\{z_{1},z_{2},\ldots,z_{n}\}$, where $z_{i}=1$ indicates a true (blocked) obstacle with underlying blockage probability $\rho^{*}_{i}$ such that $P(Z_{i}=1)=\rho^{*}_{i}$.
We denote the transformed log-odds vector as $Y=(y_{1},y_{2},\ldots,y_{n})$, where each $y_{i}=\log\left(\frac{\rho^{*}_{i}}{1-\rho^{*}_{i}}\right)$, for all i=1,2,,ni=1,2,...,n. and assume a Gaussian prior:
$Y|X\sim\mathcal{N}(\mu,K),$
where $\mu$ is the prior mean vector and $K$ is the covariance matrix. We set $\mu=0$ which centers the prior log-odds for each obstacle at zero, corresponding to a prior median (and mean, in the absence of uncertainty) of $\rho_{i}=0.5$ for all $i$. This setup indicates that each obstacle is equally likely to be true or false a priori. The matrix $K$ contains elements $\{K_{ij}=k(x_{i},x_{j})\}$ determined a kernel function $k$ capturing the inter-dependency between obstacles at two locations $x_{i}$ and $x_{j}$. We use a squared exponential kernel in spatial distance to model correlation such that nearby obstacles tend to have similar log-odds values, thereby capturing the intuition that spatially close obstacles tend to have similar blockage status:
[…]
where $\sigma_{f}$ is and $l$ are kernel parameters controlling the strength and decay rate of spatial correlations.
Given a sequence of sensor readings and disambiguation outcomes for each obstacle, denoted as $\tilde{\rho_{i}}^{1},\ldots,\tilde{\rho_{i}}^{t}$, we aim to refine the probability estimates $P(Z_{i}=1|\tilde{\rho_{i}}^{1},...,\tilde{\rho_{i}}^{t}),i=1,2,...,n$. This framework allows us to update belief over obstacles that has not been directly sensed due to limited sensing range, by using the spatial correlation captured in the Gaussian prior.
4.2 Observations
In most works addressing traversal problems in regions with potential blockages, the sensor is assumed to be noisy and conditionally dependent on the true nature of the blockage. The generated probabilistic readings about the blockage status are typically modeled using a $Beta(\alpha,\beta)$ distribution (Priebe et al., (2005), Ye et al., (2011), Aksakalli and Ari, (2014)) which has support matching the probability range.
Let $z_{i}\in\{0,1\}$ denote the true (but unobserved) status of obstacle $i$. Given $z_{i}$, the sensor reading $\tilde{\rho}_{i}$ is drawn from:
$\tilde{\rho}_{i}\mid z_{i}\sim\begin{cases}\text{Beta}(\alpha_{O},\beta_{O}),&\text{if }z_{i}=1\\ \text{Beta}(\alpha_{F},\beta_{F}),&\text{if }z_{i}=0\end{cases}$
The parameters are chosen to satisfy $\mathbf{E}(\tilde{\rho_{i}}|z_{i}=1)=\frac{\alpha_{O}}{\alpha_{O}+\beta_{O}}>\mathbf{E}(\tilde{\rho_{i}}|z_{i}=0)=\frac{\alpha_{F}}{\alpha_{F}+\beta_{F}}$, reflecting that the sensor is more likely to return higher probability marks for true obstacles and lower values for false ones, with uncertainty captured by the distribution shape parameters.
Given a new sensor reading $\tilde{\rho}_{i}^{t}$, we update its log-odds via Bayes’ rule. Assuming conditional independence of sensor readings given the obstacle statuses, the odds ratio becomes:
$\frac{P(z_{i}=1\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t})}{P(z_{i}=0\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t})}=\frac{f(\tilde{\rho}_{i}^{t}\mid z_{i}=1)P(z_{i}=1\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t-1})}{f(\tilde{\rho}_{i}^{t}\mid z_{i}=0)P(z_{i}=0\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t-1})},$
we then obtain log-odds value by taking log-transformation:
$\log\left(\frac{P(z_{i}=1\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t})}{P(z_{i}=0\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t})}\right)=\log\left(\frac{f(\tilde{\rho}_{i}^{t}\mid z_{i}=1)}{f(\tilde{\rho}_{i}^{t}\mid z_{i}=0)}\right)+\log\left(\frac{P(z_{i}=1\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t-1})}{P(z_{i}=0\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t-1})}\right).$
The resulting $\tilde{y}_{i}^{t}=\log\left(\frac{P(z_{i}=1\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t})}{P(z_{i}=0\mid\tilde{\rho}_{i}^{1},...,\tilde{\rho}_{i}^{t})}\right)$ is treated as a noisy observation of the latent true log-odds $y_{i}$, approximated by:
$\tilde{y}_{i}^{t}\approx y_{i}+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma_{i}^{2}),$
where $\sigma_{i}^{2}$ reflects the uncertainty due to the sensor noise and the informativeness of the reading $\tilde{\rho}_{i}^{t}$. Then $\tilde{y}_{i}$ serves as the input observation for the Bayesian updating process using correlation information.
To accommodate the fact that only a subset of obstacles are observed at each time step (due to limited sensing range), we assign an observation noise variance of $\sigma_{i}^{2}=\infty$ for unobserved obstacles. This effectively removes their influence from the likelihood function, ensuring that their posterior estimates are driven solely by spatial correlation with observed obstacles.
4.3 Posterior Distribution
The prior distribution on log-odds value $Y=\left(y_{1},y_{2},...,y_{n}\right)^{T}$ is assumed as:
$f(\mathbf{y})\propto\exp\left\{-\frac{1}{2}\mathbf{y}^{T}K^{-1}\mathbf{y}\right\}.$
Let $\tilde{\mathbf{y}}=(\tilde{y}_{1},\dots,\tilde{y}_{n})^{T}$ be the observed log-odds vector with independent Gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$. Then the likelihood is:
$L(\tilde{\mathbf{y}}|\mathbf{y})\propto\prod_{i=1}^{n}\exp\left\{-\frac{1}{2}\frac{\left(\tilde{y_{i}}-y_{i}\right)^{2}}{\sigma_{i}^{2}}\right\}=\exp\left\{-\frac{1}{2}\left(\tilde{\mathbf{y}}-\mathbf{y}\right)^{T}\Sigma^{-1}\left(\tilde{\mathbf{y}}-\mathbf{y}\right)\right\},$
where $\Sigma=\text{diag}(\sigma_{1}^{2},...,\sigma_{n}^{2})$. The posterior of $Y$ given observations $\tilde{Y}$ becomes a standard Bayesian posterior of multivariate normal with diagonal Gaussian noise:
$f(\textbf{y}|\tilde{\textbf{y}})\propto f(\mathbf{y})L(\tilde{\mathbf{y}}|\mathbf{y}),$
with the posterior mean and covariance:
$\mu_{pos}=\mathbf{E}\left(Y|\tilde{Y}\right)=\left[\Sigma^{-1}+K^{-1}\right]^{-1}\Sigma^{-1}\tilde{\mathbf{y}}$,
$K_{pos}=\left[K^{-1}+\Sigma^{-1}\right]^{-1}.$
The posterior mean $\mu_{pos}$ can be transformed elementwise into posterior probabilities via the logistic function:
$\rho_{i}=\frac{\exp(\mu_{pos,i})}{1+\exp(\mu_{pos,i})},\quad i=1,\dots,n.$
4.4 Benefits of Posterior Updating based on the Correlation Structure
Having established the correlation model and updating framework, we analyze the theoretical benefits of (i) incorporating spatial correlation into belief updating process and (ii) exploration through collecting new observations.
Theorem 4.1.
Let $\pi^{*}_{\mathrm{cor}}$ and $\pi^{*}_{\mathrm{ind}}$ denote the optimal traversal policies under posterior distributions with and without correlation information, respectively. Then, under the same environment,
$\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}(S_{t},d^{\pi^{*}_{\mathrm{cor}}}(S_{t}))\right]\;\leq\;\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}(S_{t},d^{\pi^{*}_{\mathrm{ind}}}(S_{t}))\right].$
Proof.
Let $\Phi_{\mathrm{cor}}$ denote the random variable representing posterior distributions over obstacle statuses when correlation information is incorporated (i.e., under the full Gaussian Random Field model). Similarly, let $\Phi_{\mathrm{ind}}$ denote the random variable when the posterior distribution is obtained under the assumption that obstacle statuses are independent.
Define a transformation $\Gamma$ that removes the correlation structure from any realization $\phi_{\mathrm{cor}}$ of $\Phi_{\mathrm{cor}}$, yielding a corresponding independent posterior:
$\Phi_{\mathrm{ind}}=\Gamma(\Phi_{\mathrm{cor}}).$
In the sense of Blackwell informativeness (Blackwell,, 1953; de Oliveira,, 2018), $\Phi_{\mathrm{ind}}$ is a garbling of $\Phi_{\mathrm{cor}}$, meaning that $\Phi_{\mathrm{cor}}$ is at least as informative as $\Phi_{\mathrm{ind}}$.
For any policy $\pi$ that maps belief states to decisions, let $\mathcal{C}(S_{t},d^{\pi}(S_{t}))$ denote the cost incurred at time $t$ under state $S_{t}$ and decision $d^{\pi}(S_{t})$. Then, by Blackwell’s Theorem, the optimal expected cost under the more informative signal (posterior with correlation) is no greater than that under the less informative signal:
$\min_{\pi\in\Pi_{\Phi_{\mathrm{cor}}}}\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi}(S_{t})\bigr)\right]\;\leq\;\min_{\pi\in\Pi_{\Phi_{\mathrm{ind}}}}\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi}(S_{t})\bigr)\right],$
where $\Pi_{\Phi_{\mathrm{cor}}}$ and $\Pi_{\Phi_{\mathrm{ind}}}$ denote the respective sets of feasible policies conditioned on their posterior models.
By definition of the optimal policies $\pi^{*}_{\mathrm{cor}}$ and $\pi^{*}_{\mathrm{ind}}$, the inequality becomes
$\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi^{*}_{\mathrm{cor}}}(S_{t})\bigr)\right]\;\leq\;\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi^{*}_{\mathrm{ind}}}(S_{t})\bigr)\right],$
which completes the proof. ∎
This result implies that correlation-aware updating provides theoretical guarantees for improved performance. We also examine how additional information monotonically improves decision quality, further supporting the goal of uncertainty reduction in our posterior beliefs by sequentially gathering new observations through sensing or disambiguation.
Corollary 4.2 (Monotonicity with additional observations).
Let $X^{U}$ be the set of currently ambiguous obstacles and let $X^{U}_{1}\subseteq X^{U}_{2}\subseteq X^{U}$ be two observation sets, and $\Phi_{1}$, $\Phi_{2}$ denote the random variables representing the posterior distributions following observing those two sets, respectively. Let $\Pi_{1}$ and $\Pi_{2}$ be the sets of feasible policies with respect to their posterior models. Then the minimum expected total cost is monotonically decreasing in additional observations:
$\mathbf{E}\left[\mathcal{C}(X^{U}_{1})\right]\geq\mathbf{E}\left[\mathcal{C}(X^{U}_{2})\right],$
where $\mathbf{E}\left[\mathcal{C}(X)\right]:=\min_{\pi\in\Pi_{X}}\mathbf{E}\!\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi}(S_{t})\bigr)\right].$ Equivalently, for the cost reduction function defined as $f(X):=\mathbf{E}\left[\mathcal{C}(\emptyset)\right]-\mathbf{E}\left[\mathcal{C}(X)\right]$, we have that $f$ is monotonically increasing: $f(X^{U}_{2})\geq f(X^{U}_{1})$.
Proof.
Let $\tilde{Y}_{1}$ and $\tilde{Y}_{2}$ denote the random log-odds vectors from sets $X^{U}_{1}$ and $X^{U}_{2}$ respectively, where $\tilde{Y}_{2}=\left(\tilde{Y}_{1},\tilde{Y}_{2\setminus 1}\right)$ and $\tilde{Y}_{2\setminus 1}$ represents the additional log-odds observations from $X^{U}_{2}\setminus X^{U}_{1}$.
For any realization of the observations:
$p(Y|\tilde{Y}_{1}=\tilde{\mathbf{y}}_{1})$ $=\int p(Y|\tilde{Y}_{1}=\tilde{\mathbf{y}}_{1},\tilde{Y}_{2\setminus 1}=\tilde{\mathbf{y}})\cdot p(\tilde{Y}_{2\setminus 1}=\tilde{\mathbf{y}}|\tilde{Y}_{1}=\tilde{\mathbf{y}}_{1})\,d\tilde{\mathbf{y}}$
$=\int p(Y|\tilde{Y}_{2}=(\tilde{\mathbf{y}}_{1},\tilde{\mathbf{y}}))\cdot p(\tilde{Y}_{2\setminus 1}=\tilde{\mathbf{y}}|\tilde{Y}_{1}=\tilde{\mathbf{y}}_{1})\,d\tilde{\mathbf{y}}$
This shows that for any realization, the posterior $\Phi_{1}$ is obtained by averaging the posterior $\Phi_{2}$ over the additional observations $\tilde{Y}_{2\setminus 1}$. Hence $\Phi_{1}$ is a garbling of $\Phi_{2}$.
By Blackwell’s Theorem (as in Theorem 4.1), the more informative signal $\Phi_{2}$ yields no greater minimum expected cost than $\Phi_{1}$, establishing the inequality $f(X^{U}_{2})\geq f(X^{U}_{1})$. ∎
5 Two-Stage Policy Learning Framework
Based on the sequential model proposed in Section 3.1, our goal is to find an optimal policy $\pi^{*}$ that guides the agent in making decisions based on the current state information. Each decision step determines not only the immediate path before encountering a new obstacle, but also impacts all future decision steps. Expanding the objective function, the optimal policy $\pi^{*}$ selects the decision $d_{t}$ from the decision set $\mathcal{D}_{t}(S_{t})$ at each step according to the following equation:
$d_{t}^{*}=\operatorname*{arg\,min}_{d_{t}\in\mathcal{D}_{t}(S_{t})}\left\{\mathcal{C}(S_{t},d_{t})+\mathbf{E}_{W_{t+1}|S_{t},d_{t}}\left[\min_{\{d_{t^{\prime}},t^{\prime}=t+1,...,T\}}\mathbf{E}\left(\sum_{t^{\prime}=t+1}^{T}\mathcal{C}(S_{t^{\prime}},d_{t^{\prime}})\Big|S_{t+1}\right)\right]\right\},$
This is a full lookahead policy (Powell,, 2022) that optimizes over the entire time horizon until reaching the goal $g$. However, we need to enumerate all possible future state-decision sequences, $\{S_{t^{\prime}},d_{t^{\prime}},W_{t^{\prime}+1}\}$, which is infeasible in practice.
Rollout policies offer a popular and computationally efficient approximation for real-time decision making (Bertsekas,, 2021). However, they are highly sensitive to underlying approximation policy quality and environmental information accuracy, making their performance less robust in highly uncertain environments (Eyerich et al.,, 2010; Hou and Srinivasa,, 2022; Blumenthal and Shani,, 2023).
To address this challenge, we propose a two-stage policy learning framework (Figure 3) for efficient decision making: (i) the offline phase learns a good estimation of the minimum expected value, $\min\mathbf{E}\left(\sum_{t^{\prime}=t+1}^{T}\mathcal{C}(S_{t^{\prime}},d_{t^{\prime}})|S_{t+1}\right)$, which induces a base policy $\pi^{\text{base}}$, and (ii) the online phase generates rollout policy $\pi^{\text{roll}}$, which uses the following rule for real-time decision making, $d_{t}^{\text{roll}}=\operatorname*{arg\,min}_{d_{t}\in\mathcal{D}_{t}(S_{t})}\left(\mathcal{C}(S_{t},d_{t})+\mathbf{E}_{W_{t+1}|S_{t},d_{t}}\mathbf{E}\left(\sum_{t^{\prime}=t+1}^{T}\mathcal{C}(S_{t^{\prime}},d_{t^{\prime}}^{\text{base}})|S_{t+1}\right)\right)$. Throughout the traversal process, we periodically update the base policy using observations from rollout phase.
[…]
Figure 3: Two-stage policy learning for SCOS
In this section, we first introduce three key components that enable our framework for effective and efficient policy learning. we then detail the offline and online policy learning in Sections 6 and 7.
5.1 Identifying the Decision Set with Search Space Reduction
At each decision step, we first need to identify the decision set $\mathcal{D}_{t}(S_{t})$, which contains the stopping vertices of all feasible macro-path segments. These stopping vertices are either vertices adjacent to ambiguous obstacles or the goal $g$. Constructing $\mathcal{D}_{t}$ at each decision step using an exhaustive search by examining all obstacles and adjacent vertices quickly becomes computationally prohibitive as the number of obstacles increases. To overcome this challenge, we introduce a heuristic approach that: (i) identifies a set of decisions associated with prioritized obstacles using an optimistic greedy search, and (ii) reduces the search space by discarding paths unlikely to be optimal.
5.1.1 Macro-Path and Decision Set Construction
Each decision $d_{t}\in\mathcal{D}_{t}(S_{t})$ moves the agent from $v_{t}$ to a location $v_{t+1}$ via a macro-path segment. Following the sensing and disambiguation (if needed), the state transitions from $S_{t}$ to $S_{t+1}$. The key question is: which ambiguous obstacles should be prioritized for disambiguation and sensing, and at which adjacent vertices (referred to as the disambiguation vertex) should this be implemented? We use an optimistic greedy approach that iteratively identifies candidates to be included in $\mathcal{D}_{t}(S_{t})$.
[…]
Input: Current state $S_{t}=(\mathcal{V}_{t},B_{t})$, goal vertex $g$
Output: Decision set $\mathcal{D}_{t}$
1 Initialize: $\mathcal{D}_{t}\leftarrow\emptyset$, $\tilde{X}^{F}\leftarrow X^{F}\cup X^{U}$, $\tilde{X}^{O}\leftarrow X^{O}$;
2
3repeat
 /* Path Selection */
4   Find minimum-cost path $p^{*}$ from $v_{t}$ to $g$ under $\{\tilde{X}^{F},\tilde{X}^{O}\}$;
5 if no ambiguous obstacle intersects $p^{*}$ then
6    $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t}\cup\{g\}$;
7    break;
8
9 else
10      Identify the first ambiguous obstacle intersected $x_{i}\in X^{U}$;
11      Let $v(x_{i})$ be the disambiguation vertex just before $x_{i}$;
12
    /* Decision Set Update */
13    $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t}\cup\{v(x_{i})\}$;
14    $\tilde{X}^{F}\leftarrow\tilde{X}^{F}\setminus\{x_{i}\}$, $\tilde{X}^{O}\leftarrow\tilde{X}^{O}\cup\{x_{i}\}$;
15
16   end if
17
18until $g\in\mathcal{D}_{t}$;
Algorithm 1 Identification of Decision Set
While including more candidates provides more information to enhance the decision-making process, it may lead to over-exploration and waste of computational resources (MacDonald and Smith,, 2020). The following proposition (Proposition 5.1) ensures that our approach (Algorithm 1) effectively avoids over-exploration and stops when no further improvement can be achieved over a pure exploitation decision (i.e., the best path avoiding all ambiguous and true obstacles based on current knowledge). Define a candidate in the decision set $\mathcal{D}_{t}$ to be a valid candidate if it can potentially achieve improvement over pure exploitation decision, we show that:
Proposition 5.1.
The decision set $\mathcal{D}_{t}$ determined by the iterative optimistic greedy approach (Algorithm 1) contains only valid candidates, either disambiguation vertices or the goal $g$.
Proof.
At each iteration, we find the minimum cost path under optimism assumption and identify the first ambiguous obstacle $x_{i}$ along this path. The cost of this path can be decomposed as:
$\ell_{free}(v_{t},x_{i})+c(x_{i})+\mathcal{C}_{optimism}(x_{i},g),$
where $\ell_{free}(v_{t},x_{i})$ represents the length of obstacle-free path from current vertex $v_{t}$ to the obstacle $x_{i}$, $c(x_{i})$ is obstacle’s disambiguation cost and $\mathcal{C}_{optimism}(x_{i},g)$ is the minimum total traversal cost from obstacle $x$ to $g$ by assuming all obstacles of uncertain status are false (i.e., traversable). Since the algorithm stops when $g$ is added to the set, it is equivalent to adding disambiguation vertices of all obstacles satisfying the following condition,
$\ell_{free}(v_{t},g)\geq\ell_{free}(v_{t},x)+c(x)+\mathcal{C}_{optimism}(x,g).$
This indicates that disambiguating $x$ could potentially yield a cheaper path than the best obstacle-free path (i.e., the pure exploitation path). Otherwise, $x$ will not be added to $\mathcal{D}_{t}$ if its best possible performance cannot beat the best obstacle-free path. ∎
5.1.2 Search Space Reduction
To further improve efficiency, we reduce the search space by discarding obstacles and their associated paths whose best performance cannot exceeding the current best solution. Combining the decision set construction with search space reduction strategy, we minimize computation cost while preserving only the most promising traversal options. This enables the optimal policy to scale effectively to larger and more complex environments.
To establish the condition for discarding each obstacle, we compare two values: (i) the upper bound of the optimal path and (ii) the lower bound of the best path reaching the goal via the obstacle. If the lower bound exceeds the upper bound, then no path via this obstacle can beat the current best path, indicating the obstacle and its associated paths can be excluded from consideration.
We first retain all obstacles identified by Algorithm 1, and the final shortest obstacle-free path identified can provide an upper bound on the optimal path cost. The reduction algorithm (Algorithm 2) then starts with evaluating the remaining obstacles. For each remaining obstacle, we calculate a lower bound on the path cost passing through it. Since each obstacle contains multiple interior vertices, finding the best path through this obstacle theoretically leads to the exhaustive search. Instead, we introduce a pseudo vertex to facilitate the calculation. Let $\mathcal{V}_{\text{in}}=\{v\in\mathcal{V}:v\in\text{interior}(x)\}$ for obstacle $x$, where the interior is approximated using a radius $\text{radius}(x)$ (e.g., derived from geometric or sensor range assumptions). A pseudo-vertex $\tilde{v}$ is then introduced, connected to each $v\in\mathcal{V}_{\text{in}}$ via zero-cost edges. The lower bound becomes the concatenation of the minimum cost obstacle-free path from current vertex to this pseudo vertex and the minimum cost path under optimism assumption from this pseudo vertex to $g$. Then obstacles whose lower bound exceeds the upper bound and the associated paths are discarded.
Input: State $S_{t}=(\mathcal{V}_{t},B_{t})$, goal vertex $g$, decision set $\mathcal{D}_{t}$, associated obstacles $X(\mathcal{D}_{t})$, upper bound $\mathcal{C}_{U}$
Output: Discarded obstacle set $X_{\text{discard}}$
1 Initialize: $X_{\text{discard}}\leftarrow\emptyset$, $\tilde{X}^{F}\leftarrow X^{F}\cup X^{U}$, $\tilde{X}^{O}\leftarrow X^{O}$;
2
3foreach $x\in X\setminus(X^{O}\cup X(\mathcal{D}_{t}))$ do
4   Identify interior vertices: $\mathcal{V}_{\text{in}}\leftarrow\{v\in\mathcal{V}:\|v-x\|\leq\text{radius}(x)\}$;
5   Introduce pseudo-vertex $\tilde{v}$ and connect to all $v\in\mathcal{V}_{\text{in}}$ via zero-cost edges;
6
 /* Compute lower bound cost via obstacle $x$ */
7 $\mathcal{C}_{1}\leftarrow\min_{p\in\mathcal{P}_{v_{t},\tilde{v}}:p\cap X^{U}=\emptyset}\ell_{p}$;
8 $\mathcal{C}_{2}\leftarrow\min_{p\in\mathcal{P}_{\tilde{v},g}}(\ell_{p}+\mathcal{C}_{p})$;
9
10 if $\mathcal{C}_{1}+\mathcal{C}_{2}\geq\mathcal{C}_{U}$ then
11    $X_{\text{discard}}\leftarrow X_{\text{discard}}\cup\{x\}$;
12
13   end if
14
15 end foreach
Algorithm 2 Search Space Reduction
How Algorithm 2 Prunes the Search Space. Algorithm 2 further reduces the planning complexity by eliminating obstacles that provably cannot contribute to lower-cost traversal paths. These obstacles are excluded from future disambiguation, simulation, and policy considerations, thereby reducing the graph size, candidate action set, and overall computational burden. This pruning step is especially valuable in the SCOS framework, where the number of ambiguous obstacles can be large, each disambiguation is costly, and evaluating traversal paths under correlated uncertainty is computationally intensive. While Algorithm 1 identifies the most promising obstacles for immediate disambiguation, Algorithm 2 complements this by removing the remaining obstacles that offer no potential benefit, resulting in a streamlined and more tractable decision process.
5.2 Information-Guided Exploration Strategy
As the agent traverses the environment, collecting new information through sensor readings and disambiguation is crucial for making decisions. However, not all information is equally valuable, and we need to quantify the information value to guide more informed exploration. We build on the mutual information (Chaloner and Verdinelli,, 1995) proposed for Bayesian experimental design problem. In the context of observations with Gaussian noise, the information gain can be calculated using (Srinivas et al.,, 2010):
$\textbf{I}(\tilde{Y}^{t})=\frac{1}{2}\log\left|I+\sigma^{-2}\Sigma_{v}\right|,$
where $\tilde{Y}^{t}$ is the random vector corresponding to potential log-odds observations can be collected, $K_{v}$ is the associated covariance matrix and $\sigma^{2}$ is the noise variance. This quantity measures how much uncertainty is reduced by incorporating new observations into the decision-making process.
In most cases, gaining information is more critical during early stages of the traversal, and as the agent approaches the goal $g$, the influence of new information decreases, which is also justified by the submodularity of information gain (see Proposition 6.2). To reflect this decreasing influence, we adapt the method from Contal et al., (2014) to compute the information gain $G(d)$ at time step $t$ of implementing any decision $d$ as:
$G(d)=\sqrt{\gamma}\left(\sqrt{\textbf{I}(\tilde{Y}^{t})+\sum_{i=1}^{t-1}\textbf{I}(\tilde{Y}^{i})}-\sqrt{\sum_{i=1}^{t-1}\textbf{I}(\tilde{Y}^{i})}\right),$
where $\gamma$ is the scaling factor controlling the weight of information gain in decision-making. We set $\gamma$ proportionally to the empirical standard deviation of the estimated value function across the decision set $\mathcal{D}_{t}$, ensuring information gain and value estimates are on similar scale.
5.3 Posterior Sampling of Environment Models
In our problem setting, the uncertainty related to the transition function (i.e., obstacle blockage statuses) requires additional step to address its associated variability throughout the navigation. Unlike traditional threshold-based methods that categorize obstacles using fixed cut-off probabilities (Koenig and Likhachev,, 2002; Bnaya et al.,, 2009) or prior sampling approaches that assume either static probabilistic information or a small subset of environment status as a prior, we use posterior sampling that evolves with accumulated observations. This approach provides environment models for both offline base policy learning and online decision making stages, offering several advantages over existing methods: (i) it automatically balances exploration and exploitation as uncertainty changes over time, (ii) it preserves spatial correlation information throughout the sampling process, and (iii) it eliminates the need for manual threshold tuning.
Specifically, using the posterior distribution derived as in Section 4, we sample log-odds, which are modeled as latent variables with correlation structure, from $f(\mathbf{y}|\tilde{\rho}^{1},...,\tilde{\rho}^{t})$ at time $t$. These log-odds are then transformed into probabilities using $\rho_{i}=\frac{\exp(y_{i})}{1+\exp(y_{i})}$. The obstacle status for each location is determined using binary sampling according to probabilities $\rho_{i}$. And $\tilde{X}=\tilde{X}^{O}\cup\tilde{X}^{F}$ is updated accordingly and helps to form the predicted traversal region $\tilde{\mathcal{G}}$ for interaction simulation. The process is outlined in Algorithm 3.
Input: $f(\mathbf{y}|\tilde{\rho}^{1},...,\tilde{\rho}^{t})$
Output: Predicted traversal region $\tilde{\mathcal{G}}$
1 Sample $(y_{1},y_{2},...,y_{n})\sim f(\mathbf{y}|\tilde{\rho}^{1},...,\tilde{\rho}^{t})$;
2 Transform $\rho_{i}=\frac{\exp(y_{i})}{1+\exp(y_{i})}$, i=1,2,,ni=1,2,...,n;
3 Sample $z_{i}\sim\text{Bernoulli}(\rho_{i})$ for each i=1,2,…,n, where $z_{i}=1$ indicates a true obstacle;
4 Update $\tilde{X}^{O}=\{x_{i}:z_{i}=1\}$, $\tilde{X}^{F}=\{x_{i}:z_{i}=0\}$;
Update traversal region $\tilde{\mathcal{G}}$ using $\tilde{X}=\tilde{X}^{O}\cup\tilde{X}^{F}$
Algorithm 3 Posterior Sampling for Environment Model Generation
Together, these three components enable our two-stage policy learning framework to develop robust base policies efficiently, and to maintain an online rollout policy that continuously adapts to updated beliefs.
6 Offline Base Policy Learning Guided by Information
The offline stage aims to learn a high-quality base policy $\pi^{\text{base}}$ that provides estimates of the expected future traversal cost following state $S_{t+1}$, denoted as the optimal state value function $V^{*}(S_{t+1})$:
$V^{*}(S_{t+1})=\min_{\{d_{t^{\prime}},t^{\prime}=t+1,...,T\}}\mathbf{E}\left(\sum_{t^{\prime}=t+1}^{T}\mathcal{C}(S_{t^{\prime}},d_{t^{\prime}})|S_{t+1}\right)\approx\mathbf{E}\left(\sum_{t^{\prime}=t+1}^{T}\mathcal{C}(S_{t^{\prime}},d^{\text{base}}(S_{t^{\prime}}))|S_{t+1}\right)$
We propose an optimistic policy iteration (OPI) framework augmented with information bonus with three key features: (i) quantifying information gain to encourage targeted exploration, (ii) integrating model-based posterior sampling with model-free value learning process, and (iii) supporting both point estimate and distribution learning approaches for stronger uncertainty quantification.
6.1 Optimistic Policy Iteration with Information Integration
Policy iteration (PI) is a foundamental framework in RL (Sutton and Barto,, 2018). Its core idea is the alternating implementation between two key steps: policy evaluation and policy improvement. The policy evaluation step involves generating interaction experience under a fixed policy. The resulting observations are used to update the state value functions. In the policy improvement step, the updated value function estimates are used to define a better policy, supporting further refinement. This alternating process drives the estimated value functions $V(s)$ towards their optimal values $V^{*}(s)$.
Unlike traditional PI that requires full convergence of value function estimates before proceeding to the policy improvement, our base policy learning uses the optimistic policy iteration (OPI) framework that allows partial updates, making it more computationally efficient and adaptive. Applying it in our problem setting, we perform a single round of value function updates before policy improvement. The key innovation in our approach is the direct incorporation of information gain built upon mutual information, and modify the Bellman optimality equation into:
$V^{*}(s)=\min_{d\in\mathcal{D}(s)}\mathcal{C}(s,d)+\mathbf{E}\left[V^{*}(s^{\prime})|s,d\right]-G(d),$
where $s^{\prime}$ represents the resulting state after implementing $d$ at state $s$. Subtracting $G(d)$ effectively rewards decisions that provide valuable information, encouraging exploration of high-information paths during the learning process. The theoretical advantage is established in Section 6.4. For notational simplicity, in all subsequent sections we denote $\mathcal{C}^{\scriptscriptstyle IG}(s,d)$ as the cost after adjusting for $G(d)$.
6.2 Model-Based and Model-Free Reinforcement Learning Structure
Within this information-guided OPI framework, we integrate model-based and model-free components. During policy evaluation, the model-based component uses posterior sampling to generate probabilistic environment models that guide exploration. The model-free component complements this by directly learning value functions from simulated experiences, refining estimates while adapting to sequentially updated posterior distributions. The interplay of these two components creates a robust learning environment, effectively addressing uncertainty in obstacle status (i.e., blockage probabilities).
Following value function updates, an improved policy minimizing expected traversal cost is defined in the policy improvement step, completing one iteration. This process repeats until convergence. We adopt a partial convergence approach focusing on stabilizing value estimates only for states in current decision set, rather than requiring full convergence across the entire state space. This approach is similar to value iteration with restricted backups, a technique commonly used in approximate dynamic programming. It not only accelerates the overall convergence of the process, but also ensures that value estimates for non-converged states serve as effective initializations for the subsequent online rollout phase.
Our approach aligns with Dyna-based methods in hybrid-RL literature (Sutton,, 1991; Silver et al.,, 2008), where the model-based component provides dynamic information to enhance model-free learning and improve sampling efficiency. By incorporating posterior sampling through Bayesian updating, we emphasize a systematic approach for handling uncertainty and optimizing exploration-exploitation trade-off, resulting in more robust base policy learning. The steps are detailed in Algorithm 4.
[…]
Output: Estimates of optimal value function $V^{*}(s)$
1 Initialize: $i\leftarrow 1$;
2while not converged do
 /* Policy Evaluation */
[…]
4   Simulate experience under $\pi_{i}$ and update $V_{i+1}(s)$ using $\mathcal{C}^{\scriptscriptstyle IG}(s,d)$;
5
 /* Policy Improvement */
6   Define $\pi_{i+1}$ using $V_{i+1}(s)$;
7
 /* Convergence Check */
8   Compute the difference between $V_{i+1}(s)$ and the updated value function $V_{i}(s)$;
9 if $\max_{s}|V_{i+1}(s)-V_{i}(s)|<\eta$ then
10     Return $V^{*}(s)\leftarrow V_{i+1}(s)$
11   end if
12
13 end while
Algorithm 4 Information-Guided OPI for Base Policy Learning
Convergence Property:
Our problem is an undiscounted, finite horizon problem, and prior work has established convergence results for OPI using single round value function updates per iteration under static probabilistic information for such problems (Tsitsiklis, (2002), Chen, (2018), Winnicki and Srikant, (2023)). However, these results do not directly cover our proposed method due to the incorporation of posterior sampling. In Section 6.4, we prove that the convergence can still be established in a similar way, complementing with extensive empirical results presented in Section 8.
6.3 The Model-Free Component: Value Function Approximation
The model-free component of our offline learning process directly estimates the optimal state value functions from simulated experience. We explore two distinct strategies to generate robust and efficient value estimation: (i) Monte Carlo approach for point estimates and (ii) distributional reinforcement learning (distributional RL) approach for estimating full distributions, capturing not only the mean but also the uncertainty.
6.3.1 Monte Carlo Approximation
The Monte Carlo (MC) approach estimates value function by using cumulative observed costs from simulated trajectories. In a sampled traversal environment, with a given policy, a complete trajectory is simulated from a starting state randomly selected from decision set $\mathcal{D}(S)$. Denote the cumulative traversal costs with information adjustment from any state $s$ on this trajectory in $i^{th}$ iteration as $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$, the value function is updated using the following rule:
$V_{i}(s)=(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}\mathcal{C}^{\scriptscriptstyle IG}_{i}(s),\quad\text{with }\alpha_{i}=\frac{1}{N_{i}(s)}$ (2)
where $N_{i}(s)$ represents the total number of times state $s$ has been experienced in all simulations up to $i^{th}$ iteration. For efficiency, we simultaneously update the value function for all states experienced along the trajectory.
After updating the state value function, we obtain a better policy in the policy improvement step to move towards the optimal direction. Combining with the Monte Carlo method, we consider three strategies for comparison.
Greedy method.
The greedy method is most commonly used, which always selects the decision that minimizes expected cost based on current value estimates:
$d_{\text{greedy}}=\arg\min_{d\in\mathcal{D}(s)}\left\{\mathcal{C}^{\scriptscriptstyle IG}(s,d)+V_{i}(s^{\prime})\right\}.$
Decaying $\epsilon$-greedy method.
Due to the uncertainty related to single trajectory simulation and transition probability sampling, we consider the $\epsilon$-greedy strategy. It introduces controlled exploration by selecting random decisions with probability $\epsilon$ while following the greedy option with probability $1-\epsilon$:
[…]
Specifically, we gradually decrease $\epsilon$ after each decision step to reflect that exploration is more critical during early stages of the traversal.
Softmax exploration method.
Although decaying $\epsilon$-greedy strategy encourages exploration, it treats all non-greedy decisions equally. Hence, we consider the softmax strategy using Boltzmann distribution, assigning the selection probability proportional to the updated value function. This strategy maintains relative preferences among decisions as below:
$P(d\mid s)=\frac{\exp\left(-\beta\left(\mathcal{C}^{\scriptscriptstyle IG}(s,d)+V_{i}(s^{\prime})\right)\right)}{\sum_{d^{\prime}\in\mathcal{D}(s)}\exp\left(-\beta\left(\mathcal{C}^{\scriptscriptstyle IG}(s,d^{\prime})+V_{i}(s^{\prime})\right)\right)},$
where we set $\beta=1$ as a default parameter.
The process of using Monte Carlo approach for estimating value function and updating policy is outlined in Algorithm 5.
Input: Traversal region $\tilde{\mathcal{G}}$, policy $\pi_{i}$, value estimates $V_{i}(s)$ for all $s$
Output: Improved policy $\pi_{i+1}$, updated value estimates $V_{i+1}(s)$
/* Trajectory Simulation */
1 Start from a random state $s_{0}\in\mathcal{D}$;
2 while trajectory not terminated do
3   Take action $d$ according to $\pi_{i}$ and observe $\mathcal{C}^{\scriptscriptstyle IG}(s,d)$, next state $s^{\prime}$;
4 $s\leftarrow s^{\prime}$;
5
6 end while
/* Value Update */
7 foreach state $s$ along trajectory do
8   Accumulate $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ for each visited $s$;
9 $V_{i+1}(s)\leftarrow(1-\alpha)V_{i}(s)+\alpha\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$;
10
11 end foreach
/* Policy Improvement */
12 if selection is greedy then
13 $\pi_{i+1}(s)\leftarrow\arg\min_{d\in\mathcal{D}(s)}\,\bigl\{\mathcal{C}^{\scriptscriptstyle IG}(s,d)+V_{i+1}(s^{\prime})\bigr\}$.
14 else if selection is $\epsilon$-greedy then
15 $\pi_{i+1}(s)$ chooses a random $d$ w.p. $\epsilon$, else the greedy decision.
16 else if selection is softmax then
17 $\pi_{i+1}(s)$ selects $d$ w.p. proportional to $\exp\bigl(-\beta[\mathcal{C}^{\scriptscriptstyle IG}(s,d)+V_{i+1}(s^{\prime})]\bigr)$.
Algorithm 5 Monte Carlo approach for value function and policy update
6.3.2 Distributional RL Approach
While Monte-Carlo approach generates point estimates (i.e., expected values) of $V(s)$, it does not capture the inherent uncertainty in estimations arising from transition and posterior sampling. To handle this uncertainty, we propose a distributional RL learning approach, a more robust method capturing the full distributions of $V(s)$. Moreover, the exploration strategies paired with point estimate methods often require careful exploration parameter tuning, which may lead to inconsistent performance compared to distribution-based approaches like posterior sampling (Osband et al.,, 2013; Osband and Van Roy,, 2017).
Using the distributional Bellman equation, Bellemare et al., (2017) introduced a distributional reinforcement learning (distributional RL) algorithm using categorical distribution, called C51 algorithm. This algorithm approximates the discrete distribution of $V(s)$, demonstrating significant improvements over point estimate approaches. Expanding on this foundation, Dabney et al., (2018) proposed Quantile Regression DQN (QR-DQN), which utilizes quantile regression to represent continuous distributions, enhancing the flexibility of value function representation. Focusing on addressing exploration challenges, Mavrin et al., (2019) designed an exploration bonus using quantile-based distributional RL techniques, while Tang and Agrawal, (2018) unified posterior sampling with distributional RL to create an exploration framework, shown to improve exploration performance in policy learning.
Given that we have a finite number of obstacle statuses in the environment, the value function $V(s)$ can only take a finite number of possible values. Therefore, we adopt the categorical distribution to represent the probability distribution over $V(s)$. Extending the approach of Tang and Agrawal, (2018), we propose a framework using Bayesian updates to learn the distribution of $V(s)$. Posterior sampling from sequentially updated distributions drives the policy improvement, which is demonstrated to be more stable and efficient compared to the classic exploration strategies described in the previous section. The complete process is outlined in Algorithm 6.
Bayesian update.
Let $V(s)$ be a random variable following a categorical distribution over finite support set $\mathcal{X}=\{x_{1},\dots,x_{k}\}$:
$V(s)\sim\text{Cat}(\mathbf{p})\quad\text{where }\mathbf{p}=(p_{1},\dots,p_{k}),$
and $p_{i}=p(V(s)=x_{i}),i=1,2,...,k$ represents the probability that $V(s)$ takes the value $x_{i}\in\mathcal{X}$. We place a Dirichlet prior over probabilities:
$\mathbf{p}=\left(p_{1},p_{2},...,p_{k}\right)\sim\text{Dir}(\bm{\alpha}=\left(\alpha_{1},\alpha_{2},...\alpha_{k}\right)),$
where $\alpha_{1}=\alpha_{2}=\dots=\alpha_{k}=1$ corresponds to a uniform prior, indicating no preference among the possible values. Given observed costs from state $s$ over the first $i$ iterations, $\{\mathcal{C}^{\scriptscriptstyle IG}_{1}(s),\mathcal{C}^{\scriptscriptstyle IG}_{2}(s),...,\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)\}$, the posterior of $\mathbf{p}$ is derived using Bayes’ rule:
$f(\mathbf{p}|\{\mathcal{C}^{\scriptscriptstyle IG}_{1}(s),\mathcal{C}^{\scriptscriptstyle IG}_{2}(s),...,\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)\})\propto f(\mathbf{p};\bm{\alpha})\prod_{i}f(\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)|\mathbf{p}).$
By conjugacy, the resulting posterior remains Dirichlet:
$\mathbf{p}\mid\{\mathcal{C}^{\scriptscriptstyle IG}_{1}(s),\dots,\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)\}\sim\text{Dir}(\bm{\alpha}^{\prime}=\left(\alpha^{\prime}_{1},\alpha^{\prime}_{2},...\alpha^{\prime}_{k}\right)),$ (3)
where $\alpha^{\prime}_{k}=\alpha_{k}+\sum\mathbbm{1}\{\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)=x_{k}\}$ with $\mathbbm{1}\{\cdot\}$ being the indicator function.
Recursive update using distributional Bellman equation.
Following a policy $\pi$ at state $s$, we observe immediate information-adjusted cost $\mathcal{C}^{\scriptscriptstyle IG}(s,d)$ and next state $s^{\prime}$. The core idea of distributional RL relies on the distributional Bellman equation of random variables $V(s)$:
$V^{\pi}(s)=_{D}\mathcal{C}^{\scriptscriptstyle IG}(s,d)+V^{\pi}(s^{\prime}),$
where $d$ is determined by policy $\pi$. Based on the distribution of $V^{\pi}(s^{\prime})$, let $\mathcal{X}(V(s^{\prime}))=\{x^{\prime}_{1},x^{\prime}_{2},...x^{\prime}_{k^{\prime}}\}$ denote the support of $V(s^{\prime})$, we update $V(s)$ by treating $\mathcal{C}^{\scriptscriptstyle IG}(s,d)+x^{\prime}_{i}$, $i=1,2,\dots,k^{\prime}$ as observations. However, the shifted values $\mathcal{C}^{\scriptscriptstyle IG}(s,d)+\mathcal{X}(V(s^{\prime}))$ often does not align with the support $\mathcal{X}(V(s))$. Following Bellemare et al., (2017), we resolve this discrepancy using weighted interpolation. For each $x^{\prime}_{i}\in\mathcal{X}(V(s^{\prime}))$ with associated probability $p^{\prime}_{i}$, we compute the shifted value $x^{\prime}_{i}+\mathcal{C}^{\scriptscriptstyle IG}(s,d)$, which falls in the interval $x_{j},x_{j+1}]$ (i.e., $x_{j}\leq x^{\prime}_{i}+\mathcal{C}^{\scriptscriptstyle IG}(s,d)<x_{j+1}$). Then using Bayes updating rule we derived, we update the Dirichlet parameters for $V(s)$:
$\alpha_{j}=\frac{|x_{j+1}-(x^{\prime}_{i}+\mathcal{C}^{\scriptscriptstyle IG}(s,d))|}{x_{j+1}-x_{j}}p^{\prime}_{i}+\alpha_{j},\;\alpha_{j+1}=\frac{|x_{j}-(x^{\prime}_{i}+\mathcal{C}^{\scriptscriptstyle IG}(s,d))|}{x_{j+1}-x_{j}}p^{\prime}_{i}+\alpha_{j+1}.$
Support refinement through value contraction.
To initialize the distribution of $V(s)$, we determine their upper and lower bounds using the method described in Section 5.1, then discretize the interval into an equally spaced grid with spacing $\delta$ (i.e., $\mathcal{X}=\{x_{\text{L}},x_{\text{L}}+\delta,x_{\text{L}}+2\delta,\dots,x_{\text{U}}\}$). While practical and commonly used in distributional RL algorithms, this fixed discretization method does not guarantee accurate representation of the true value distribution. To address this shortcoming, we propose a support refinement step using a value contraction approach, which adds observed values or replaces the unobserved ones.
During the simulation, the distribution of $V(s)$ is updated iteratively using the distributional Bellman equation at each step. After completing an interaction trajectory, we refine the support of $V(s)$ using the observed information-adjusted costs. Let $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ represent the cumulative information-adjusted cost from state $s$ observed during the $i^{th}$ iteration, we update the distribution support by locating neighbors of $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ in $\mathcal{X}(V(s))$. Specifically, we identify two consecutive support points $x_{j},x_{j+1}$ (i.e., $x_{j}\leq\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)\leq x_{j+1}$) such that
$|x_{j}-\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)|\leq\delta,\;|x_{j+1}-\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)|\leq\delta.$
If these neighboring points have not been observed in previous trajectories, we replace them with $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ and update its probability distribution parameter:
$\alpha_{j}=\alpha_{j}+\alpha_{j+1}+1.$
Otherwise, we directly insert $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ into the support set and assign $\alpha_{j+1}=1$, since it is the first observation of that value. If $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ is already included in the support, we update using the Bayesian update rule. This refinement guarantees that the support of $V(s)$ aligns with the actual values which can be observed, enhancing the accuracy of distribution representation.
Input: Environment $\tilde{\mathcal{G}}$, initial policy $\pi_{i}$, value distribution $V_{i}(s)$ for all $s$
Output: Improved policy $\pi_{i+1}$, updated value distribution $V_{i+1}(s)$
/* Trajectory Simulation */
1 Start from a random state $s_{0}\in\mathcal{D}$;
2while trajectory not terminated do
3   Take action $d\sim\pi_{i}(s)$, Observe $\mathcal{C}^{\scriptscriptstyle IG}(s,d)$ and next state $s^{\prime}$;
4
 /* Distributional Bellman Update */
5   Compute distribution $V_{i+1}(s)\leftarrow\mathcal{C}^{\scriptscriptstyle IG}(s,d)+V_{i}(s^{\prime})$;
6   Update $V_{i}(s)$ using weighted interpolation: $\alpha^{\prime}_{j}=\frac{|x_{j+1}-(x^{\prime}_{i}+\mathcal{C}^{\scriptscriptstyle IG}(s,d))|}{x_{j+1}-x_{j}}p^{\prime}_{i}+\alpha_{j},\;\alpha^{\prime}_{j+1}=\frac{|x_{j}-(x^{\prime}_{i}+\mathcal{C}^{\scriptscriptstyle IG}(s,d))|}{x_{j+1}-x_{j}}p^{\prime}_{i}+\alpha_{j+1}$;
7
8 $s\leftarrow s^{\prime}$;
9
10 end while
11
/* Support Refinement */
12 foreach state $s$ along trajectory do
13   Compute cumulative cost $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)=\sum_{t=0}^{T}\mathcal{C}^{\scriptscriptstyle IG}(s_{t},d_{t})$;
14   Identify neighbors $x_{j},x_{j+1}$ in $\mathcal{X}(V_{i+1}(s))$: $|x_{j}-\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)|\leq\delta,\;|x_{j+1}-\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)|\leq\delta$;
15 if $x_{j},x_{j+1}$ not observed and $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)\notin\mathcal{X}(V_{i}(s))$ then
16      Replace $x_{j},x_{j+1}$ with $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$;
17      Update $\alpha_{j}\leftarrow\alpha_{j}+\alpha_{j+1}+1$;
18
19 else
20    if $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)\in\mathcal{X}(V_{i}(s))$ then
21         Update $\alpha_{j}$ using Equation (3);
22
23    else
24       $\alpha_{j}\leftarrow 1$
25      end if
26
27   end if
28
29 end foreach
30
/* Policy Improvement */
31 foreach state $s$ do
32   Sample $\mathbf{p}$ from posterior distribution and compute $\mathbf{E}\left[V(s)\right]$;
33 $\pi_{i+1}\leftarrow\operatorname*{arg\,min}_{d}\mathcal{C}^{\scriptscriptstyle IG}(s,d)+\mathbf{E}\left[V(s)\right]$
34 end foreach
35
Algorithm 6 Distributional RL with Bayesian update and support refinement for value function and policy update
6.4 Offline Learning Framework Properties
In this section, we establish theoretical results to justify our offline learning strategy. First, we demonstrate the benefit of incorporating information gain into value function updating process, then show that the information gain should maintain diminishing impact on the decision making due to submodularity property. We also prove the convergence of OPI under posterior sampling.
Theorem 6.1.
Given that the state space and the decision set are finite in our problem setting, assume every policy $\pi$ is proper (i.e., reaches the goal with probability 1 in finitely many steps), and the immediate cost $\mathcal{C}(s,d)$ is bounded, $\sup_{s,d}\mathcal{C}(s,d)<\infty$. Define two Bellman operators
$(\mathbf{T}V)(s)=\min_{d\in\mathcal{D}(s)}\bigl\{\mathcal{C}(s,d)+\mathbf{E}[V(s^{\prime})\mid s,d]-G(d)\bigr\},\;(\mathbf{T}_{\mathrm{b}}V)(s)=\min_{d\in\mathcal{D}(s)}\bigl\{\mathcal{C}(s,d)+\mathbf{E}[V(s^{\prime})\mid s,d]\bigr\}.\$
Let $V^{\!*}$ and $V^{b}$ denote their fixed points. Then:
  1. (i)
    $V^{*}(s)\leq V^{b}(s)$ for all $s\in\mathcal{S}$.
  2. (ii)
    Let $\pi^{b}$ be the optimal policy obtained from $V^{b}$. Then for any start state $s_{0}$,
    $V^{b}(s_{0})-V^{*}(s_{0})\geq\mathbf{E}_{\pi^{b}}\left[\sum_{t=0}^{T-1}G(d_{t})\right],$
    where the expectation is taken over trajectory distribution induced by $\pi^{b}$ and transition probabilities.
Proof.
Because $G(d)\geq 0$, we have point-wise operator ordering $(\mathbf{T}V)(s)\leq(\mathbf{T}_{\mathrm{b}}V)(s)$ for every bounded $V$ and state $s$. Bellman operators are monotone contractions, hence iteration from any common initial value preserves this order, leading to $V^{\!*}\leq V^{b}$.
Given a state $s$, denote
$d^{b}(s)=\pi^{b}(s),\quad\Delta(s)=V^{b}(s)-V^{\!*}(s)\;\;\geq 0.$
Because $V^{\!*}$ is the minimum in $\mathbf{T}$, it satisfies, for any decisionădd,
$V^{\!*}(s)\leq\mathcal{C}(s,d)+\mathbf{E}[V^{\!*}(s^{\prime})\mid s,d]-G(d),$
where $s^{\prime}$ is the next random state given current state-decision pair. Let $d=d^{b}$ and subtracting the equation for $V^{b}$ results in the inequality:
$\Delta(s)\;\geq\;G(d^{b})\;+\;\mathbf{E}\bigl[\Delta(s^{\prime})\mid s,d^{b}\bigr].$
Cumulatively, it leads to
$\Delta(s_{0})\;\geq\;\sum_{t=0}^{T-1}G(d^{b}_{t}),$
taking expectation with respect to the trajectory distributions induced by $\pi^{b}$ and transition probabilities yields the desired result. ∎
Proposition 6.2 (Submodularity of information gain).
Consider the log-odds vector $Y\sim\mathcal{N}(0,K)$ in our SCOS model. Let $\tilde{Y}_{X^{\prime}}$ be the random vector corresponding to potential log-odds observations of obstacle set $X^{\prime}\subseteq X$. In the context of observations with Gaussian noise, define the information gain as the mutual information
$\mathbf{I}(\tilde{Y}_{X^{\prime}})=\text{MI}(Y_{X^{\prime}};\tilde{Y}_{X^{\prime}})=\tfrac{1}{2}\log\det\!\big(I+\Sigma_{X^{\prime}}^{-1}K_{X^{\prime}X^{\prime}}\big),$
where $K_{X^{\prime}X^{\prime}}$ is the covariance matrix corresponding to the obstacle set $X^{\prime}$, and $\Sigma_{X^{\prime}}$ is the diagonal matrix containing observation noise variance. Then:
  1. (i)
    Normalization: $\mathbf{I}(\varnothing)=0$.
  2. (ii)
    Monotonicity: if $X_{1}\subseteq X_{2}\subseteq X$, then $\mathbf{I}(\tilde{Y}_{X_{2}})\geq\mathbf{I}(\tilde{Y}_{X_{1}})$.
  3. (iii)
    Submodularity: for all $X_{1}\subseteq X_{2}\subseteq X$ and any $x_{i}\notin X_{2}$,
    $\mathbf{I}(\tilde{Y}_{X_{1}\cup\{x_{i}\}})-\mathbf{I}(\tilde{Y}_{X_{1}})\;\geq\;\mathbf{I}(\tilde{Y}_{X_{2}\cup\{x_{i}\}})-\mathbf{I}(\tilde{Y}_{X_{2}}).$
Proof.
Normalization. $\mathbf{I}(\varnothing)=\tfrac{1}{2}\log\det(I)=0$.
Monotonicity. Let $X_{2}=X_{1}\cup\{x_{i}\}$. By the chain rule for mutual information,
$\mathbf{I}(\tilde{Y}_{X_{2}})-\mathbf{I}(\tilde{Y}_{X_{1}})=\text{MI}\big(Y_{x_{i}};\tilde{Y}_{x_{i}}\mid\tilde{Y}_{X_{1}}\big).$
For observations with Gaussian noise,
$\text{MI}\big(Y_{x_{i}};\tilde{Y}_{x_{i}}\mid\tilde{Y}_{X_{1}}\big)=\tfrac{1}{2}\log\Big(1+\tfrac{\text{Var}(Y_{x_{i}}\mid\tilde{Y}_{X_{1}})}{\sigma_{i}^{2}}\Big)\geq 0.$
So, $\mathbf{I}(\tilde{Y}_{X_{2}})\geq\mathbf{I}(\tilde{Y}_{X_{1}})$.
Submodularity. For $X_{1}\subseteq X_{2}$ and $x_{i}\notin X_{2}$,
$\mathbf{I}(\tilde{Y}_{X_{1}\cup\{x_{i}\}})-\mathbf{I}(\tilde{Y}_{X_{1}})=\tfrac{1}{2}\log\Big(1+\tfrac{\text{Var}(Y_{x_{i}}\mid\tilde{Y}_{X_{1}})}{\sigma_{i}^{2}}\Big).$
We can calculate $\mathbf{I}(\tilde{Y}_{X_{2}\cup\{x_{i}\}})-\mathbf{I}(\tilde{Y}_{X_{2}})$ in a similar way. In a Gaussian model, conditioning on more observations can only reduce conditional variance, $\text{Var}(Y_{x_{i}}\mid\tilde{Y}_{X_{2}})\leq\text{Var}(Y_{x_{i}}\mid\tilde{Y}_{X_{1}})$. Since $f(u)=\tfrac{1}{2}\log(1+u/\sigma_{i}^{2})$ is increasing, the marginal gain after observing $X_{1}$ is at least as large as after observing $X_{2}$, proving submodularity. ∎
Remark 6.3 (Justification for greedy sensing and disambiguation).
Because $\textbf{I}(\cdot)$ is normalized, monotone, and submodular, the standard greedy selection achieves a (11/e)(1-1/e) approximation for maximizing I under a cardinality or budget constraint. This justifies our use of a simple greedy sensing and disambiguation strategy.
Theorem 6.4 (Convergence under posterior sampling).
Given state space and decision set are finite in our problem setting, assume every policy $\pi$ is proper, and the immediate cost is bounded $\sup_{s,d}|\mathcal{C}^{\scriptscriptstyle IG}(s,d)|<\infty$. Let $\bar{\bm{\rho}}$ be the mean transition probabilities derived from the posterior distribution over log-odds value $f(\mathbf{y}|\tilde{\rho}^{1},...,\tilde{\rho}^{t})$. The sequence $\{V_{i}\}$ generated by OPI framework with posterior sampling according to the update rule in Equation (2) converges almost surely to $V^{*}$ under transition information $\bar{\bm{\rho}}$.
Proof.
Define the Bellman operator under transition model $\bm{\rho}$ for a fixed policy $\pi$ as
$(\mathbf{T}_{\pi}^{\bm{\rho}}V)(s)=\mathcal{C}^{\scriptscriptstyle IG}(s,d^{\pi}(s))+\mathbf{E}_{\bm{\rho}}\left[V(s^{\prime})\right],$
and the optimal Bellman operator as
$(\mathbf{T}^{\bm{\rho}}V)(s)=\min_{d}\left\{\mathcal{C}^{\scriptscriptstyle IG}(s,d(s))+\mathbf{E}_{\bm{\rho}}\left[V(s^{\prime})\right]\right\}.$
Let the update rule be $V_{i}(s)=(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$, where $\alpha_{i}=\frac{1}{N_{i}(s)}$, satisfying $\sum\alpha_{i}=\infty$ and $\sum\alpha_{i}^{2}<\infty$. $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ can be decomposed as $V^{\pi_{i},\bm{\rho}^{i}}+\omega_{i}$, where $V^{\pi_{i},\bm{\rho}^{i}}=\mathbf{T}_{\pi_{i}}^{\bm{\rho}^{i}}V_{i-1}$, and $\omega_{i}$ is the zero mean noise. Rewriting this update rule as
$V_{i}(s)=(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}V^{\pi_{i},\bm{\rho}^{i}}+\alpha_{i}\omega_{i}.$
Define the posterior sampling noise as
$\eta_{i}=\mathbf{T}_{\pi_{i}}^{\bm{\rho}^{i}}V_{i-1}-\mathbf{T}_{\pi_{i}}^{\bar{\bm{\rho}}}V_{i-1},$
which satisfies $\mathbf{E}\left[\eta_{i}\right]=0$ due to the linearity of expectation. Following Lemma 2.5 in Chen, (2018), we have:
$\mathbf{T}_{\pi_{i}}^{\bar{\bm{\rho}}}V_{i-1}\leq\mathbf{T}^{\bar{\bm{\rho}}}V_{i-1}+\bm{\epsilon}.$
Since $\mathbf{E}\left[\omega_{i}\right]=0$, the inequality is established:
$V_{i}(s)$ $=(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}V^{\pi_{i},\bm{\rho}^{i}}+\alpha_{i}\omega_{i}$
$\leq(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}\bigl[\mathbf{T}^{\bar{\bm{\rho}}}V_{i-1}+\bm{\epsilon}\bigr]+\alpha_{i}(\eta_{i}+\omega_{i})$
$=(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}\,\mathbf{T}^{\bar{\bm{\rho}}}V_{i-1}+\alpha_{i}\,\bm{\epsilon}+\alpha_{i}\,\tilde{\omega}_{i},$
where $\tilde{\omega}_{i}(s)$ is the combined zero-mean noise term. Then following the arguments in the proof of Theorem 2.1 in Chen, (2018) and Proposition 1 in Tsitsiklis, (2002), we have
$\lim_{i\rightarrow\infty}V_{i}=V^{*}\;\text{almost surely}.$
7 Online Rollout with Base Policy Update
Building on the high-quality base policy learned offline, our online phase implements rollout policy for making real-time decisions. Rollout policies are theoretically guaranteed to perform no worse than their underlying base policy (Bertsekas,, 2021), making the quality of the base policy crucial for overall performance. Formally, the rollout policy makes decisions following the rule:
$d_{t}^{\text{roll}}=\operatorname*{arg\,min}_{d_{t}\in\mathcal{D}_{t}(S_{t})}\left(\mathcal{C}(S_{t},d_{t})+\mathbf{E}_{W_{t+1}|S_{t},d_{t}}\left[V^{\text{base}}(S_{t+1})\right]\right),$
where $V^{\text{base}}(S_{t+1})=\min\mathbf{E}\left[\sum_{t^{\prime}=t+1}^{T}\mathcal{C}(S_{t^{\prime}},d_{t^{\prime}}^{\text{base}})|S_{t+1}\right]$ is the estimated optimal state value function following the base policy.
At $t=0$, the decision can be made directly using the pre-computed $V^{\text{base}}$ from offline stage. From $t=1$ onward, simulation is required to incorporate the updated belief as new information is collected. Given the performance guarantee of rollout policies, maintaining base policy quality becomes critical as belief information evolves.
Base Policy Update Mechanism
A unique challenge in the SCOS setting is that the belief over obstacle blockage statuses is not static. Instead, it evolves as new information is collected through sensor readings and disambiguations. If left unadjusted, $\pi^{\text{base}}$ may become misaligned with the improved belief, decreasing the effectiveness of the rollout policy. To address this challenge, we update $\pi^{\text{base}}$ after each rollout step, ensuring that our two-stage framework remains robust and adaptive throughout the traversal process.
The updating process varies depending on the value function approximation method used in the offline phase. Under point estimation approach, we update $V^{\text{base}}$ using the incremental learning rule from Equation (2). Under the distributional RL approach, we update the value distribution using the support refinement process outlined in Section 6.3.2.
8 Monte Carlo Experiments and Comparison
We evaluate policies using 8-adjacency grid graphs that align with the discretized setting of the real-world COBRA minefield dataset, which is commonly used in path planning studies (Priebe et al., (2005), Ye et al., (2011), Aksakalli et al., (2016), Aslan et al., (2020)).
Formally, the traversal region is a two-dimensional plane, $\Omega=[0,I]\times[0,J]$, represented by an undirected 8-adjacency integer lattice graph $\mathcal{G}$. $\mathcal{V(G)}$ is the set of all vertices corresponding to grid intersections, and $\mathcal{E(G)}$ is the set of edges connecting each vertex to its horizontal, vertical and diagonal neighbors, corresponding to the sides and diagonals of the grid squares. Each vertex $v\in\mathcal{V(G)}$ has a pair of integer coordinates $(i,j)$, where i=1,2,,Ii=1,2,...,I and j=1,2,,Jj=1,2,...,J. Each edge $e\in\mathcal{E(G)}$ is defined between vertices of the following forms: $(i,j)$ and $(i+1,j)$, $(i,j)$ and $(i,j+1)$, $(i,j)$ and $(i+1,j+1)$. We set the start vertex at $s=(\lfloor I/2\rfloor,J)$ and the goal vertex as $g=(\lfloor I/2\rfloor,1)$, encouraging strategic navigation around obstacles rather than detouring along an unnecessarily long obstacle-free path. Grid sizes range from compact (i.e, 50×25) to larger area operations (i.e, 100×50), reflecting different problem scales.
Disk-shaped obstacles of fixed radius are randomly positioned within a subregion of $\Omega$ to ensure that at least one long traversable path exits between $s$ and $g$, with obstacle density in the traversal region span from lightly defended (i.e, 20-30 obstacles) to heavily filled (i.e, 40-60 obstacles). To reflect realistic minefield characteristics, we impose a spatial correlation pattern between obstacles based on two practical defensive strategies: (i) obstacles closer to the goal has higher log-odds of being blocked (i.e., true threats), simulating defensive strategy for protecting objectives, and (ii) isolated obstacles (with fewer neighboring obstacles) have higher log-odds compared to clustered ones, reflecting surveillance patterns where decoy clusters hide real threat locations. The resulting setup ensures obstacles with similar tactical importance exhibit correlated log-odds values, with additional noise generated according to multivariate normal distribution for capturing environmental uncertainties.
The sensor readings are generated from $Beta(\alpha,\beta)$, where parameters vary depending on sensor noise and the actual obstacle status:
$\alpha_{O}=4+\lambda,\;\beta_{O}=4-\lambda,\;\text{if }z_{i}=1$
$\alpha_{F}=4-\lambda,\;\beta_{F}=4+\lambda,\;\text{if }z_{i}=0$
$\lambda\in(0,4)$ controls the level of sensor noise, with lower values (e.g., $\lambda=0.35,0.75$) simulating basic detectors with high false alarm rates, and higher values (e.g., $\lambda=1.5,2.5$) representing advanced radar with improved discrimination capabilities. Sensor ranges vary from restricted capabilities (i.e., 10-15 grid units) to extended detection ranges (i.e., 20-30 grid units), reflecting different operational constraints.
Graph Size Number of Obstacles ($N$)
Disambiguation Cost
(= Radius)
Sensor Range ($R$)
(to obstacle center)
Sensor Parameter ($\lambda$)
$50\times 25$ 2020 $3.5$ $10,12.5,15$ $0.35,0.75,1.5,2.5$
$40$ 33
$100\times 50$ $30$ $5.5$ $20,25,30$ $0.35,0.75,1.5,2.5$
$60$ 55
Table 1: Parameters of Simulation Setting
The complete simulation parameter combinations are summarized in Table 1. For each combination, we generate 50 distinct environments, each includes 10 random replicates according to the probability information, yielding 500 total simulation runs per parameter setting. Example traversal regions of different obstacle density levels are presented in Figure 4.
Figure 4: Example traversal environments containing $N=40$ and $N=20$ obstacles (red, gray disks present obstacles that are actual threats and false alarms, respectively)
8.1 Evaluation Metrics
To comprehensively compare policy performance, we use the following evaluation metrics:
Average performance.
We compute the mean and median traversal costs across all simulation environments, providing direct comparison of policy efficiency and typical performance.
Optimality gap.
We assess each policy’s deviation from optimality by calculating the difference between achieved costs and an offline-optimal benchmark: $\mathcal{C}-\mathcal{C}_{\text{optimal}}$, quantifying how closely each policy approaches the optimal solution. This offline benchmark assumes perfect knowledge of obstacle actual status, which is an unrealistic assumption in practice.
Consistency and robustness.
We evaluate (i) within-environment consistency by calculating the standard deviation of traversal costs across replicates within each environment, and (ii) cross-environment robustness by calculating the standard deviation of mean traversal costs across environments, assessing policy reliability under stochastic conditions.
Convergence speed.
We record the average offline and online simulation time per complete traversal, indicating the policy appropriateness for real-time decision making.
8.2 Competing Policies
We consider four competing baseline approaches which represent diverse strategies for solving navigation problems, ranging from computationally efficient heuristics to approaches effectively incorporating probabilistic information, leading to a balanced evaluation of our proposed framework.
Penalty-based policies.
These approaches assign a deterministic value to each path by penalizing high-risk ones in addition to the actual traversal length, then apply classic shortest path algorithms (e.g., Dijkstra’s algorithm) with replanning when encountering an ambiguous obstacle. We consider two variants (Sahin and Aksakalli,, 2015; Alkaya et al.,, 2021): (i) RD policy penalizes cost of path using $\tilde{\mathcal{C}}_{p}=\ell_{p}+\sum_{x:x\cap p\neq\emptyset}\frac{c_{x}}{1-\rho_{x}}$, and (ii) DT policy incorporates distance-to-goal term $d(x,g)$, shown to have comparable performance to UCT-based methods: $\tilde{\mathcal{C}}_{p}=\ell_{p}+\sum_{x:x\cap p\neq\emptyset}\left[c_{x}+\left(\frac{d(x,g)}{1-\rho_{x}}\right)^{-\log(1-\rho_{x})}\right]$.
Rollout-based policies.
These approaches use simulation to evaluate future trajectories, but differ in the base policy strategies and assumptions compared to our two-stage learning framework. While we learn high-quality base policy offline and continuously adapt them to evolving beliefs, these baselines rely on simple heuristic assumptions that remain static throughout traversal. We consider two approaches (Eyerich et al.,, 2010; Hou and Srinivasa,, 2022): (i) hindsight policy using rollout with a base policy that assumes perfect information about sampled environment during simulation, serving as a powerful benchmark for comparison in the literature, and (ii) optimistic rollout policy assuming all ambiguous obstacles in sampled environment are traverasable (i.e., optimistic assumption) during simulation, encouraging exploration in uncertain regions.
8.3 Illustrative Example
We use one example traversal region including 40 potentially blocked disks to illustrate how different policies and correlation modeling impact decisions and traversal cost.
Figure 5(a) presents the traversal from our proposed two-stage policy learning framework. The path takes a short, direct route after the information collection through sensing and disambiguation in the region near the starting vertex, reaching the goal with total cost 46.21. Figure 5(b)-5(c) include the traversal results using the baseline, DT policy, under two modeling assumptions, correlation-aware and independent. Compared with our two-stage strategy, the penalty policy is conservative and leads to much higher traversal cost. With correlation modeling, DT policy treats the central clustered region as high risk and tends to take detour around the region, resulting in a route with cost 90.36. The performance is worse when correlation is ignored. The policy cannot refine beliefs on statuses of nearby obstacles, repeatedly disambiguates locally with additional costs, increasing the cost to 120.43.
(a) Two-Stage Policy Learning
(b) Penalty Policy (correlation)
(c) Penalty Policy (independent)
Figure 5: Illustration using region with 40 potentially blocked disks under different policies and correlation assumptions.
8.4 Empirical Results
We present key numerical results and insights in this section, with additional results for different combinations of environmental parameters included in the Appendix.
Figure 6 displays the mean traversal costs with 95% confidence intervals across environments. In general, traversal costs decrease monotonically as $\lambda$ increases (i.e., sensor precision increases), graph size reduces and obstacle density decreases, confirming the expected relationship between environmental complexity and path planning difficulty. Across policies, our proposed two-stage policy framework consistently yields lower mean traversal costs and tighter confidence interval than rollout and penalty baselines, especially in more challenging environments with high noise and dense obstacles, showing the sensitivity of baselines to environmental factors. Exceptions are observed when sensors are highly accurate, where the baselines appear to be comparably effective. These exceptions likely occur because high sensor accuracy condition creates a relatively straightforward setting that does not require sophisticated uncertainty handling techniques. In such simple scenarios, iterative learning approaches tend to be conservative, but performance can be improved via tuning exploration parameter (e.g., weight of information gain).
Within the two-stage framework, the distributional RL (DRL) base exhibits the lowest traversal costs in most cases, with particularly strong performance in challenging environments due to its superior uncertainty quantification. Due to better exploration mechanisms, the Monte Carlo bases using softmax and $\epsilon$-greedy exploration strategies, due to better exploration mechanisms demonstrate competitive performance and establish advantage over the greedy base. Among baseline policies, the optimistic rollout policy appears as the strongest competitor, while the hindsight policy performs surprisingly poorly. Comparing two penalty policies, the RD policy consistently exhibits high traversal costs, particularly under conditions of high obstacle density and high noise, while DT policy shows better performance across most settings.
Median cost comparisons (see Figure A1) show similar performance ranking pattern, with more pronounced advantages of using the proposed policy learning framework, further validates its benefits.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: The mean traversal cost with 95% confidence intervals for proposed policy variants and baselines
Figures A2 and A3 show the mean and median traversal costs by sensor range. The traversal cost decreases as sensor range increases, with reduction magnitude growing with sensor range. This validates the effectiveness of our Bayesian updating framework, it converts additional observations into better decisions, and the gains are further amplified by incorporating correlation information.
Beyond mean and median cost, we assess policy performance relative to the offline-optimal benchmark that assumes perfect knowledge of obstacle status. Figure 7 presents the average optimality gap (i.e., $\mathcal{C}-\mathcal{C}_{\text{optimal}}$) across environments, revealing consistent performance rankings. The two-stage framework with DRL base achieves the smallest optimality gap across most environments, with this advantage most pronounced in larger, denser and more noised environments where uncertainty management is harder. The Monte Carlo bases, decaying $\epsilon$-greedy and softmax approaches, rank following the DRL base, maintaining competitive performance. The optimistic rollout policy shows comparable performance in smaller settings, and other baselines show comparable performance only when sensor accuracy is highest. Results aggregated by sensor range shows similar pattern and are presented in the Appendix (Figure A4).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: The average deviation from optimal solutions with 95% confidence intervals for proposed policy variants and baselines
Figure 8 presents the standard deviation of traversal costs across replicates within each environment and across environments, providing insights about policy’s robustness. DRL demonstrates higher consistency with lower variance, which is particularly crucial in applications requiring predictable policy behavior. All Monte Carlo bases show comparable robustness with slightly higher variance. Among baselines, only the optimistic rollout policy maintains reasonable consistency, while other three baselines exhibit substantially higher variance, which is getting worse in more challenging environments.
Refer to caption Refer to caption
Figure 8: The standard deviation of traversal cost across replicates and environments for proposed policy variants and baselines
Table 2 summarizes the average simulation time per complete traversal. Both online simulation time and offline training time grow with obstacle density and traversal region size for all policies. Within the two-stage policy learning framework, the DRL base takes longer computation time than Monte Carlo bases, reflecting the additional computational cost associated with distributional updates, with the smaller gap in small instances. By contrast, the decay $\epsilon$-greedy and softmax bases achieve comparable traversal performance at substantially lower computation time. Their online time is often comparable to, or shorter than, rollout baselines due to the offline training effort. Rollout baselines require online simulation time on the similar scale (or slightly faster), but come with higher traversal costs and variability. Penalty policies, which use direct cost approximation without iterative learning, run extremely fast but associate with much worse and less stable performance.
Table 2: Average simulation time (in seconds) for one complete traversal
(a) Online
NN TS(greedy) TS($\varepsilon$-greedy) TS(softmax) TS(DRL) RO(optimism) RO(hindsight) penalty(RD) penalty(DT)
20 9.22 (0.65) 7.92 (0.56) 9.56 (0.67) 18.09 (1.26) 13.28 (0.77) 16.49 (1.03) 0.05 (0.01) 0.04 (0.01)
30 20.02 (2.29) 17.13 (2.13) 18.86 (2.13) 43.04 (4.92) 40.24 (5.25) 17.81 (1.72) 0.59 (0.03) 0.48 (0.04)
40 71.16 (5.70) 57.16 (5.66) 72.48 (5.12) 208.67 (15.49) 74.29 (3.28) 70.61 (4.85) 0.36 (0.02) 0.22 (0.01)
60 316.57 (17.02) 195.13 (15.27) 311.72 (15.38) 557.36 (30.04) 194.80 (10.32) 135.31 (5.05) 0.43 (0.01) 0.26 (0.01)
(b) Offline
NN TS(greedy) TS($\varepsilon$-greedy) TS(softmax) TS(DRL)
20 19.03 (0.93) 19.75 (0.95) 19.77 (0.96) 28.97 (1.35)
30 56.49 (4.38) 61.40 (4.75) 57.10 (4.41) 99.58 (7.38)
40 89.02 (3.49) 85.15 (3.37) 92.87 (3.53) 252.08 (7.75)
60 347.95 (12.86) 282.38 (11.18) 343.15 (12.49) 477.80 (17.97)
8.5 Summary of Key Empirical Findings
Bayesian updating framework.
The proposed Bayesian updating process demonstrates effectiveness in using sensor readings to improve decisions, with greater benefits under challenging conditions where uncertainty handling and information efficiency is crucial. Consistent with Theorem 4.1 and Corollary 4.2, which show that added observations reduce the expected cost and correlation-aware updating process amplify these gains.
Two-stage policy learning framework.
Across environments, the two-stage framework outperforms rollout and penalty baselines on mean, median cost, optimality gap, and standard deviations, with benefits most pronounced under more challenging conditions (high uncertainty and noise, dense obstacles). This aligned with our theoretical analysis in Section 6.4, where we show the convergence property and the exploration benefit by incorporating information gain.
Two-stage framework with distributional RL base.
DRL base appears to be the best performer, achieving the lowest mean traversal costs, smallest optimality gaps and great consistency across majority of environmental settings. This shows that its ability to handle uncertainty through distribution learning is particularly beneficial in high-noise environments where sensor information is unreliable. Its computational requirement is justified by the robust gains in accuracy and stability.
Two-stage framework with Monte Carlo bases.
While showing slightly worse performance compared to DRL, Monte Carlo bases provide a balance between performance and computational efficiency, making them suitable for settings where computational constraints are tight. With the flexibility offered by exploration parameters in decaying $\epsilon$-greedy and softmax approaches, they have the potential to achieve performance improvement through strategic tuning, which is a promising direction for future research.
Comparison baselines.
The optimistic rollout policy shows promising performance in simpler cases but shows higher costs and variability in challenging settings, limiting its practical applicability. The penalty policies provide great implementation simplicity but underperform other iterative learning approaches, making them suitable only when the computation cost is the dominant constraint or the environment is simple with accurate sensors. The hindsight policy consistently shows poor performance across all metrics and is not ideal in practical applications.
9 Discussion and Conclusions
We address a path-planning problem in complex environments of limited and uncertain information. We introduce the Stochastic Correlated Obstacle Scene (SCOS) problem, extending the Stochastic Obstacle Scene (SOS) problem by incorporating realistic obstacle spatial patterns and practical sensor constraints. To overcome limitations in existing planning policies, we propose a two-stage policy learning framework that integrates an offline training phase guided by information gain and an online decision phase. In the offline phase, we learn a robust base policy via optimistic policy iteration augmented with information bonus to encourage exploration in uncertainty regions. while efficient online rollout policy is applied for real-time decision making, followed by base policy adjustment. This framework systematically balances exploration-exploitation trade-offs and is supported by theoretical analysis.
Our contributions can be summarized across three key aspects. First, we formulate the SCOS problem as a more applicable framework for path-planing problems involving adversarial interruption and information uncertainty. Second, we develop an novel two-stage policy learning framework: offline learning enhanced with information bonus built upon mutual information for better exploration, with online rollout with periodic base updates for new information adaptation. This strategy yields robust policies with theoretical guarantees, supporting both Monte Carlo estimates and distributional RL for full distribution approximation. Third, using Gaussian random field model, we incorporate a Bayesian updating framework for information refinement, which not only enhances the decision-making, but also supports the search space reduction step to improve the computational efficiency.
Comprehensive empirical results demonstrate substantial performance improvements over existing baseline policies. In terms of solution quality, the proposed two-stage strategy achieves lower traversal costs and smaller optimality gaps across environments. The distributional RL base shows strongest performance in challenging environments with high noise and more obstacles, and its advantage grows as environmental complexity (i.e., noise and obstacle dense level) increases. The Monte Carlo bases are often close to DRL in traversal performance. Regarding computation time, empirical results show a clear trade-off between solution quality and speed. While distribution RL approach requires longer computation time, it offers consistency and robustness of performance, whereas Monte Carlo approaches provides alternatives for applications with computational constraints.
Despite these improvements, we have various directions for future research. (i) Computational scalability: the computational requirements of distribution RL approach may still limit its applicability in time critical scenarios, a more efficient distribution approximation techniques can be explored to enhance scalability. (ii) Exploration parameter tuning: the incorporation of information gain $G$ in our framework requires careful tuning of its weight in decision making to avoid being conservative and missing beneficial but uncertain paths. (iii) Dynamic environment extension: the current framework assumes static obstacle locations and status, while real-world applications often involve dynamic environments where obstacles status or locations evolve over time, posing additional planning challenge. (iv) Multi-agent extension: extending to multi-agent scenarios can greatly increase its applicability but which would require effective coordination or competing strategies between agents.
Appendix
Table A1: Description of notations used in the manuscript
Notation Description
Problem Setup
$\Omega$ Two-dimensional traversal region
$X$ Set of all obstacle locations
$X^{F}$ Set of false obstacle locations
$X^{O}$ Set of true obstacle locations
$X^{U}$ Set of ambiguous/uncertain obstacle locations
$\text{radius}(x)$ Radius of obstacle at location $x$
$\mathcal{G}$ Undirected graph imposed over $\Omega$
$\mathcal{V}(\mathcal{G})$ Set of vertices (navigation locations)
$\mathcal{E}(\mathcal{G})$ Set of feasible edges
$s,g$ Starting and goal vertices
Sensor and Beliefs
$z_{i}\in\{0,1\}$ Latent true status of obstacle $x_{i}$ ($1=$ true/blocked, $0=$ false/free)
$\rho_{i}$ Posterior blockage probability estimate for obstacle $x_{i}$
$\rho^{*}_{i}$ True underlying blockage probability for obstacle $x_{i}$
$y_{i}=\log\frac{\rho_{i}^{*}}{1-\rho_{i}^{*}}$ log-odds for obstacle $x_{i}$
$\tilde{\rho}_{i}$ (or $\tilde{y}_{i}$) Noisy observation for obstacle $x_{i}$
$R$ Sensor range
$c(x)$ Disambiguation cost for obstacle $x$
Sequential Decision Formulation
$S_{t}=\{V_{t},B_{t}\}$ State variables at time $t$
$V_{t}$ Agent’s physical location at time $t$
$B_{t}$ Belief state at time $t$
$d_{t}$ Decision at time $t$
$W_{t}$ Exogenous information at time $t$
$\mathcal{P}_{sg}$ Set of all paths from $s$ to $g$
$L_{p}$ Random variable for traversal length of path $p$
$C_{p}$ Random variable for disambiguation costs of path $p$
$\pi^{*}$ Optimal policy
$T$ Random arrival time at goal
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure A1: The median traversal cost with 95% confidence intervals for proposed policy variants and baselines
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure A2: The mean traversal cost with 95% confidence intervals for proposed policy variants and baselines by sensing ranges
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure A3: The median traversal cost with 95% confidence intervals for proposed policy variants and baselines by sensing ranges
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure A4: The average deviation from optimal solutions with 95% confidence intervals for proposed policy variants and baselines by sensing ranges